### v3

@@ Abstract
Connectivity (or equivalently, unweighted maximum flow) is an important measure in graph theory and combinatorial optimization. Given a graph $G$ with vertices $s$ and $t$, the connectivity $\lambda(s,t)$ from $s$ to $t$ is defined to be the maximum number of edge-disjoint paths from $s$ to $t$ in $G$.
Much research has gone into designing fast algorithms for computing connectivities in graphs. Previous work showed that it is possible to compute connectivities for all pairs of vertices in directed graphs with $m$ edges in $\tilde{O}(m^{\omega})$ time [Chueng, Lau, and Leung, FOCS 2011], where $\omega\in[2,2.3716)$ is the exponent of matrix multiplication. For the related problem of computing “small connectivities,” it was recently shown that for any positive integer $k$, we can compute $\min(k,\lambda(s,t))$ for all pairs of vertices $(s,t)$ in a directed graph with $n$ nodes in $\tilde{O}((kn)^{\omega})$ time [Akmal and Jin, ICALP 2023].
@@ 1 Introduction
Suppose the graph $G$ has $n$ vertices and $m$ edges. Computing an individual connectivity value in $G$ is easy: since the maximum flow between a fixed pair of nodes can be found in almost linear time [CKL+23], we know that for any fixed pair of vertices $(s,t)$ in $G$, we can compute $\lambda(s,t)$ in almost-optimal $m^{1+o(1)}$ time. For many applications however, knowing a single connectivity is not so useful, and it is instead far more informative to know the values of multiple connectivities in a graph.
This motivates the All-Pairs Connectivity (APC) problem, where we are tasked with computing connectivities for all pairs of vertices in a given graph. A long line of work recently culminated in a near-optimal $\tilde{O}(n^{2})$ time algorithm for solving APC over undirected graphs [AKL+22]. Throughout this paper, we focus on the general case where $G$ is directed.
@@
We can of course solve APC naively in $n^{2}m^{1+o(1)}$ time, simply by solving a separate instance of maximum flow for each pair of vertices. In dense graphs, this naive approach is actually the fastest known algorithm for APC! In slightly sparse graphs however, we can do better and solve APC in $\tilde{O}(m^{\omega})$ time [CLL13], where $\omega$ is the exponent of matrix multiplication (i.e., $\omega$ is the smallest positive real such that two $a\times a$ matrices can be multiplied using $a^{\omega+o(1)}$ arithmetic operations). The current fastest algorithms for matrix multiplication imply that $\omega<2.3716$ [WXXZ23]. If $\omega=2$, then the $\tilde{O}(m^{\omega})$ time algorithm is always at least as fast as the naive approach.
The lack of progress in finding faster algorithms for APC has motivated researchers to consider relaxations of APC, including the $k$-Bounded All-Pairs Connectivity ($k$-APC) problem:
@@
The $k$-APC problem is relevant in contexts where knowing the precise connectivity values between “well-connected” nodes is not important, and instead we care more about distinguishing for each pair of vertices whether its connectivity is small or large (where $k$ is our cutoff for what counts as “small” and “large”). Since the connectivity between any pair of nodes in $G$ is at most $n-1$, the $k$-APC problem recovers the general APC problem when $k=n-1$. As $k$ gets smaller, $k$-APC intuitively becomes easier. Indeed, it was recently shown that for any positive integer $k$, $k$-APC can be solved in $\tilde{O}((kn)^{\omega})$ time [AJ23].
@@
Theorem 1.1.
There is an algorithm solving APC in $\tilde{O}(m^{\omega})$ time.
Theorem 1.2.
There is an algorithm solving $k$-APC in $\tilde{O}((kn)^{\omega})$ time.
The algorithms establishing Theorems 1.1 and 1.2 are straightforward. However, the proofs of correctness for these algorithms presented in the literature are not at all obvious, and involve arguments in a somewhat complicated “flow vector framework.”
In this note, we present simpler proofs of correctness for the APC and $k$-APC algorithms from [CLL13] and [AJ23] respectively.
@@
The algorithms for APC and $k$-APC from [CLL13] and [AJ23] work in similar ways. Each algorithm first constructs a certain random matrix $M$ whose rows and columns are indexed by edges of $G$. Then, for each pair of vertices $(s,t)$, the algorithms return the value of
$\operatorname{rank}\leavevmode\nobreak\ M[E_{\text{out}}(s),E_{\text{in}}(t)]$
@@
See [AJ23, Section 3] for a more detailed overview of the approach in previous work.
@@
Finally, our generating function approach also yields somewhat more expressive results than the flow vector framework, leading to an easier proof of correctness for the $k$-APC algorithm in particular. Previously in [AJ23, Section 4], to prove correctness of the $k$-APC algorithm the authors had to manually reprove “low-rank” versions of all the flow vector framework results previously shown by [CLL13] for the general APC problem. This is not necessary in our approach: once we establish our generating function for edge-disjoint walks in $G$, some small additional reasoning yields both the APC and $k$-APC algorithms.
The only technical wrinkle in our approach is that our combinatorial view of the problem involves manipulating formal power series. However, as we discuss later in Remark 4.16, even this ingredient can be removed if one only wishes to establish Theorems 1.1 and 1.2 over directed acyclic graphs (which is already an interesting result, perhaps more suitable for teaching these algorithms in a classroom setting).
@@
In Section 2 we identify notation and assumptions used throughout the paper. In Section 3 we review standard definitions and properties of formal power series. In Section 4, we construct a matrix of formal power series whose entries enumerate families of edge-disjoint walks in a graph. In Sections 5 and 6 we leverage the construction from Section 4 to give simple proofs of Theorems 1.1 and 1.2 respectively. We conclude in Section 7 by mentioning some connections between our arguments and classical results in combinatorics and computer science, and highlighting open problems related to computing connectivities.
@@ 2 Preliminaries
We recall the following well-known results concerning matrix computation.
@@
Proofs of Propositions 2.1 and 2.2 can be found in [SCWW21] and [IMH82] respectively.
@@
For an accessible proof of Proposition 2.3, see [MR95, Theorem 7.2].
@@ 3 Power Series Preliminaries
We also note that if one is interested in solving APC and $k$-APC only in the special case of directed acyclic graphs, then it suffice to work with polynomials (no formal power series are needed) and this section can be skipped. We discuss this simplification for acyclic graphs in detail later on in Remark 4.16.
@@
Throughout this paper, we consider matrices whose entries are formal power series. Such matrices naturally arise when computing the inverses of polynomial matrices.
@@
Since every entry of $X$ has zero constant term, every nonzero entry of $X^{\ell}$ has degree at least $\ell$. Consequently, the infinite sum from the right-hand side of Equation 3 is well-defined, because for any $\boldsymbol{d}\in\mathbb{N}^{J}$, only finitely many terms contribute to the coefficient of
@@ 4 Enumeration with Edge-Adjacency Matrices
For every pair of edges $(e,f)$ in $G$ such that $\operatorname{head}(e)=\operatorname{tail}(f)$ (i.e., edge $e$ enters the vertex that edge $f$ exits), we introduce an indeterminate variable $x_{ef}$.
Let $X$ be the $m\times m$ matrix with rows and columns indexed by edges of $G$, such that for each pair of edges $(e,f)$ in $G$, we have
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
@@
$\det\leavevmode\nobreak\ \Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\pi\in\mathfrak{S}(S,T)}\sum_{\begin{subarray}{c}\\ (\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$ (9)
@@
Figure 1: Given walks $W_{i}=\langle e_{1},e_{2},g,e_{3},e_{4}\rangle$ and $W_{j}=\langle f_{1},f_{2},g,f_{3},f_{4}\rangle$ overlapping at $g$, we can swap their suffixes to produce walks $W^{\prime}_{i}=\langle e_{1},e_{2},g,f_{3},f_{4}\rangle$ and $W^{\prime}_{j}=\langle f_{1},f_{2},g,e_{3},e_{4}\rangle$ which still overlap at $g$. The weights for both pairs $\xi(W_{i},W_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})$ and $\xi(W^{\prime}_{i},W^{\prime}_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})$ agree because each pair traverses the same multiset of consecutive pairs of edges.
@@
In the proof of Lemma 4.7, our argument used the fact that we work over a field of characteristic two. This restriction on the characteristic is only included for the sake of simplicity, and is not necessary for Lemma 4.7 to hold. Indeed, if we instead worked over a field of odd characteristic, then all that changes is that terms in the expansion of the determinant come with a sign, and we can now pair up and cancel terms with opposite signs.
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$
@@
Let $S,T\subseteq E$ be subsets of edges of size $|S|=|T|=r$. If the graph $G$ contains $r$ edge-disjoint walks from $S$ to $T$, then $G$ also contains $r$ edge-disjoint paths from $S$ to $T$
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]$ (10)
given by Corollary 4.10. Moreover, any collection of paths $\mathcal{C}^{\prime}\neq\mathcal{C}$ from $S$ to $T$ has weight $\xi(\mathcal{C}^{\prime})\neq\xi(\mathcal{C})$, because $\mathcal{C}$ consists of edge-disjoint paths (so looking at the variables appearing in $\xi(\mathcal{C})$, we can recover $\mathcal{C}$ uniquely). Hence, no other term from the summation in Corollary 4.10 produces the same monomial $\xi(\mathcal{C})$. So $\xi(\mathcal{C})$ appears in Equation 10 with nonzero coefficient, which implies that the determinant from Equation 10 is a nonzero formal power series.
@@
DAGs are an interesting case for APC and $k$-APC, since we do not know of algorithms solving these problems on DAGs faster than the case of general directed graphs, and the best conditional lower bounds we have for these problems hold in the case of DAGs [AGI+19].
Focusing on the setting of DAGs leads to two simplifications.
First, any walk in a DAG has length at most $n-1$. Hence over DAGs, $X^{\ell}$ is the all-zeros matrix for $\ell\geq n$, so Proposition 3.3 shows that $\Gamma=(I-X)^{-1}$ is a matrix whose entries are polynomials. Thus, determinants of submatrices of $\Gamma$ are just polynomials, instead of formal power series involving infinite sums. Second, any walk in a DAG is a path. Thus we can skip the step in Lemma 4.12 of going from edge-disjoint walks to edge-disjoint paths.
@@ 5.1 Random Evaluation
$(I-X)^{-1}[S,T]=\frac{(\operatorname{adj}(I-X))[S,T]}{\det\leavevmode\nobreak\ (I-X)}.$ (12)
@@
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\frac{\det\leavevmode\nobreak\ (\operatorname{adj}(I-X))[S,T]}{Q^{r}}.$ (13)
@@
$Q^{r}=\left(\det\leavevmode\nobreak\ (I-X)\right)^{r}$
@@
$\lambda(s,t)=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)].$
@@
$\lambda=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$
@@
$\lambda(s,t)=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$

### v4

@@ Abstract
Connectivity (or equivalently, unweighted maximum flow) is an important measure in graph theory and combinatorial optimization. Given a graph $G$ with vertices $s$ and $t$, the connectivity $\lambda(s,t)$ from $s$ to $t$ is defined to be the maximum number of edge-disjoint paths from $s$ to $t$ in $G$. Much research has gone into designing fast algorithms for computing connectivities in graphs. Previous work showed that it is possible to compute connectivities for all pairs of vertices in directed graphs with $m$ edges in $\tilde{O}(m^{\omega})$ time [Chueng, Lau, and Leung, FOCS 2011], where $\omega\in[2,2.3716)$ is the exponent of matrix multiplication. For the related problem of computing “small connectivities,” it was recently shown that for any positive integer $k$, we can compute $\min(k,\lambda(s,t))$ for all pairs of vertices $(s,t)$ in a directed graph with $n$ nodes in $\tilde{O}((kn)^{\omega})$ time [Akmal and Jin, ICALP 2023].
@@ 1 Introduction
Suppose the graph $G$ has $n$ vertices and $m$ edges. Computing an individual connectivity value in $G$ is easy: since the maximum flow between a fixed pair of nodes can be found in almost-linear time [almost-linear-maxflow], we know that for any fixed pair of vertices $(s,t)$ in $G$, we can compute $\lambda(s,t)$ in almost-optimal $m^{1+o(1)}$ time. For many applications however, knowing a single connectivity is not so useful, and it is instead far more informative to know the values of multiple connectivities in a graph.
This motivates the All-Pairs Connectivity (APC) problem, where we are tasked with computing connectivities for all pairs of vertices in a given graph. A long line of work recently culminated in a near-optimal $\tilde{O}(n^{2})$ time algorithm for solving APC over undirected graphs [near-quadratic-gomory-hu]. Throughout this paper, we focus on the general case where $G$ is directed.
@@
We can of course solve APC naively in $n^{2}m^{1+o(1)}$ time, simply by solving a separate instance of maximum flow for each pair of vertices. In dense graphs, this naive approach is actually the fastest known algorithm for APC! In slightly sparse graphs however, we can do better and solve APC in $\tilde{O}(m^{\omega})$ time [CheungLauLeung2013], where $\omega$ is the exponent of matrix multiplication (i.e., $\omega$ is the smallest positive real such that two $a\times a$ matrices can be multiplied using $a^{\omega+o(1)}$ arithmetic operations). The current fastest algorithms for matrix multiplication imply that $\omega<2.3716$ [mmult-life]. If $\omega=2$, then the $\tilde{O}(m^{\omega})$ time algorithm is always at least as fast as the naive approach. The lack of progress in finding faster algorithms for APC has motivated researchers to consider relaxations of APC, including the $k$-Bounded All-Pairs Connectivity ($k$-APC) problem:
@@
The $k$-APC problem is relevant in contexts where knowing the precise connectivity values between “well-connected” nodes is not important, and instead we care more about distinguishing for each pair of vertices whether its connectivity is small or large (where $k$ is our cutoff for what counts as “small” and “large”). Since the connectivity between any pair of nodes in $G$ is at most $n-1$, the $k$-APC problem recovers the general APC problem when $k=n-1$. As $k$ gets smaller, $k$-APC intuitively becomes easier. Indeed, it was recently shown that for any positive integer $k$, $k$-APC can be solved in $\tilde{O}((kn)^{\omega})$ time [k-APC].
@@
{restatable}
theoremapc There is an algorithm solving APC in $\tilde{O}(m^{\omega})$ time. {restatable}theoremkapc There is an algorithm solving $k$-APC in $\tilde{O}((kn)^{\omega})$ time.
The algorithms establishing Sections 1 and 1 are straightforward. However, the proofs of correctness for these algorithms presented in the literature are not at all obvious, and involve arguments in a somewhat complicated “flow vector framework.”
In this note, we present simpler proofs of correctness for the APC and $k$-APC algorithms from [CheungLauLeung2013] and [k-APC] respectively.
@@
The algorithms for APC and $k$-APC from [CheungLauLeung2013] and [k-APC] work in similar ways. Each algorithm first constructs a certain random matrix $M$ whose rows and columns are indexed by edges of $G$. Then, for each pair of vertices $(s,t)$, the algorithms return the value of
$\operatorname{rank}~{}M[E_{\text{out}}(s),E_{\text{in}}(t)]$
@@
See [k-APC, Section 3] for a more detailed overview of the approach in previous work.
@@
Finally, our generating function approach also yields somewhat more expressive results than the flow vector framework, leading to an easier proof of correctness for the $k$-APC algorithm in particular. Previously in [k-APC, Section 4], to prove correctness of the $k$-APC algorithm the authors had to manually reprove “low-rank” versions of all the flow vector framework results shown by [CheungLauLeung2013] for the general APC problem. This is not necessary in our approach: once we establish our generating function for edge-disjoint walks in $G$, some small additional reasoning yields both the APC and $k$-APC algorithms.
The only technical wrinkle in our approach is that our combinatorial view of the problem involves manipulating formal power series. However, as we discuss later in Remark 4.16, even this ingredient can be removed if one only wishes to establish Sections 1 and 1 over directed acyclic graphs (which is already an interesting result, perhaps more suitable for teaching these algorithms in a classroom setting).
@@
In Section 2 we identify notation and assumptions used throughout the paper. In Section 3 we review standard definitions and properties of formal power series. In Section 4, we construct a matrix of formal power series whose entries enumerate families of edge-disjoint walks in a graph. In Sections 5 and 6 we leverage the construction from Section 4 to give simple proofs of Sections 1 and 1 respectively. We conclude in Section 7 by mentioning some connections between our arguments and classical results in combinatorics and computer science, and highlighting open problems related to computing connectivities.
@@ 2 Preliminaries
We recall the following results concerning matrix computation.
@@
Proofs of Propositions 2.1 and 2.2 can be found in [lecture-notes-inversion] and [matrix-rank] respectively.
@@
For an accessible proof of Proposition 2.3, see [Motwani1995, Theorem 7.2].
@@ 3 Power Series Preliminaries
We also note that if one is interested in solving APC and $k$-APC only in the special case of directed acyclic graphs, then it suffices to work with polynomials (no formal power series are needed) and this section can be skipped. We discuss this simplification for acyclic graphs in detail later on in Remark 4.16.
@@
In this paper, we work with matrices whose entries are formal power series. Such matrices naturally arise when computing inverses of polynomial matrices.
@@
Since every entry of $X$ has constant term zero, every nonzero entry of $X^{\ell}$ has degree at least $\ell$. Consequently, the infinite sum from the right-hand side of Equation 3 is well-defined, because for any $\boldsymbol{d}\in\mathbb{N}^{J}$, only finitely many terms contribute to the coefficient of
@@ 4 Enumeration with Edge-Adjacency Matrices
For every pair of edges $(e,f)$ in $G$ such that $\operatorname{head}(e)=\operatorname{tail}(f)$ (i.e., edge $e$ enters the vertex that edge $f$ exits), we introduce an indeterminate variable $x_{ef}$. Let $X$ be the $m\times m$ matrix with rows and columns indexed by edges of $G$, such that for each pair of edges $(e,f)$ in $G$, we have
@@
$\det~{}(I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
@@
$\det~{}\Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\pi\in\mathfrak{S}(S,T)}\sum_{\begin{subarray}{c}\\ (\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$ (9)
@@
Figure 1: Given walks $W_{i}=\langle e_{1},e_{2},g,e_{3},e_{4}\rangle$ and $W_{j}=\langle f_{1},f_{2},g,f_{3},f_{4}\rangle$ overlapping at $g$, we can swap their suffixes to produce walks $W^{\prime}_{i}=\langle e_{1},e_{2},g,f_{3},f_{4}\rangle$ and $W^{\prime}_{j}=\langle f_{1},f_{2},g,e_{3},e_{4}\rangle$ which still overlap at $g$. The weight $\xi(W_{i},W_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})$ of the first pair $(W_{i},W_{j})$ is precisely equal to the weight $\xi(W^{\prime}_{i},W^{\prime}_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})$ of the second pair $(W^{\prime}_{i},W^{\prime}_{j})$, because each pair traverses the same multiset of consecutive pairs of edges.
@@
In the proof of Lemma 4.7, our argument used the fact that we work over a field of characteristic two. This restriction on the characteristic is only included for the sake of simplicity, and is not necessary to enumerate families of edge-disjoint walks. If we instead worked over a field of odd characteristic, then all that changes is that terms in the expansion of the determinant from Lemma 4.5 come with a sign, and we can now pair up and cancel terms with opposite signs to prove a signed variant of Lemma 4.7.
@@
$\det~{}(I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$
@@
Let $S,T\subseteq E$ be subsets of edges of size $|S|=|T|=r$. If the graph $G$ contains $r$ edge-disjoint walks from $S$ to $T$, then $G$ also contains $r$ edge-disjoint paths from $S$ to $T$.
@@
$\det~{}(I-X)^{-1}[S,T]$ (10)
given by Corollary 4.10. Moreover, any collection of walks $\mathcal{C}^{\prime}\neq\mathcal{C}$ from $S$ to $T$ has weight $\xi(\mathcal{C}^{\prime})\neq\xi(\mathcal{C})$, because $\mathcal{C}$ consists of edge-disjoint paths (so looking at the variables appearing in $\xi(\mathcal{C})$, we can recover $\mathcal{C}$ uniquely). Hence, no other term from the summation in Corollary 4.10 produces the same monomial $\xi(\mathcal{C})$. So $\xi(\mathcal{C})$ appears in Equation 10 with nonzero coefficient, which implies that the determinant from Equation 10 is a nonzero formal power series.
@@
DAGs are an interesting case for APC and $k$-APC, since we do not know of algorithms solving these problems on DAGs faster than the case of general directed graphs, and the best conditional lower bounds we have for these problems hold in the case of DAGs [ap-bounded-mincut].
Focusing on the setting of DAGs leads to two simplifications. First, any walk in a DAG has length at most $n-1$. Hence over DAGs, $X^{\ell}$ is the all-zeros matrix for $\ell\geq n$, so Proposition 3.3 shows that $\Gamma=(I-X)^{-1}$ is a matrix whose entries are polynomials. Thus, determinants of submatrices of $\Gamma$ are just polynomials, instead of formal power series involving infinite sums. Second, any walk in a DAG is a path. Thus we can skip the step in Lemma 4.12 of going from edge-disjoint walks to edge-disjoint paths.
@@ 5.1 Random Evaluation
$(I-X)^{-1}[S,T]=\frac{(\operatorname{adj}(I-X))[S,T]}{\det~{}(I-X)}.$ (12)
@@
$\det~{}(I-X)^{-1}[S,T]=\frac{\det~{}(\operatorname{adj}(I-X))[S,T]}{Q^{r}}.$ (13)
@@
$Q^{r}=\left(\det~{}(I-X)\right)^{r}$
@@
$\lambda(s,t)=\operatorname{rank}~{}(I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)].$
@@
$\lambda=\operatorname{rank}~{}(I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$
@@
$\lambda(s,t)=\operatorname{rank}~{}(I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$
@@
1
2Compute the matrix $M=(I-A)^{-1}$.
For each pair of vertices $(s,t)$, return
$\operatorname{rank}M[E_{\text{out}}(s),E_{\text{in}}(t)]$
as the value of $\lambda(s,t)$.
Algorithm 1 The algorithm solving APC from [CheungLauLeung2013].